# When to Localize? A Risk-Constrained Reinforcement Learning Approach

### v1

@@ V-C Comparing with Baseline
We compared all algorithms in terms of the number of localization actions and success rates (that is, reaching the goal). The SP, TP, and CC-POMCP results represent the average performance of each policy type. We found that we can categorize the algorithms into two main groups: (M2xL), (M3xL), and RiskRL tend to have higher success rates but also localize more often while TP, BaseRL, and CC-POMCP tend to localize less but suffer from lower success rates. Our results showcase the strengths of the RiskRL planner. In our experiments, we set $c=0.4$ for general training and the policy successfully achieves this threshold. The success rate in the testing environments is also anticipated to be lower than the risk threshold, but not by a significant margin. Although, the RiskRL planner performs comparably worse to (M3xL) in the train environment as they both have a success rate of  93% with (M3xL) needing fewer localizations, the RiskRL planner generalized very well to the test-complex and test-simple environments. In the test-complex environment, our RiskRL planner used a similar amount of localizations to (M3xL) while having a 13% higher success rate. In the test-simple environment, our RiskRL planner achieved a 99% success rate, albeit with more localizations than (M3xL). One of the main strengths of RiskRL is that it can keep consistently high success rates, on par if not better than (M2xL) while having constantly lower number of localizations than (M2xL) and even on par number of localizations than (M3xL) in the test-complex environment.
We found that BaseRL ends up performing slightly better than the TP for both metrics in train and test-simple. However, when looking at the test-complex environment, BaseRLwas able to somewhat generalize to the environment and achieve a 35% success rate. In contrast, TP was only able to achieve a 21% success rate.
Compared to CC-POMCP, our results indicate that our BaseRL model performs very similarly with a slightly higher success rate but also a slightly higher number of localizations in the train environment. While every other planner can compute the next action nearly instantly, it takes CC-POMCP around 20 minutes to get through a trajectory.
Figure 6: (a) Training Graph showing Number of Localizations (b) Training Graph showing Success Rate. Refers to train environment. For RiskRL, $c$ was set to 0.4.
Figure 7: (a) Number of Localizations in each environment (b) Success Rate in each environment.
@@ V-D Analysis of RL Policies
Figure 8: Probability of localization based on the robot’s estimated location.
Figure 8 showcases each of the RL planners’ decision-making based on their estimated localization. In general, we see that BaseRL concentrates its localizations in fewer locations that are either close to obstacles or around a turn. On the other hand, RiskRL is much more uniform where it is willing to localize at more locations along the path with higher concentrations around obstacles.
@@ V-E Effect of Transition and Observation Noise
Figure 9: (a) Number of Localizations vs. Transition Noise Levels. (b) Success Rate vs. Transition Noise Levels. The results show training performance in train.
Varying Transition Noise. Figures 9(a) and 9(b) illustrate the number of localizations and success rates for RiskRL as we vary the transition noise. In all the experiments, the allowable probability of failure was set to $c=0.4$. Not surprisingly, the number of localization increases as the transition noise increases. But more crucially, we observe that RiskRL can respect this constraint even when the transition noise is very high, reaffirming the correctness of the algorithm.
Figure 10: (a) Number of Localizations vs. Different Observation Noise Levels. (b) Success Rate vs. Different Observation Noise Levels
Varying Observation Noise. In the previous results, we had no observation noise when the robot chose to obtain a measurement. Here, we study the effect of the observation noise level on the performance of RiskRL. Figures 10(a) and 10(b) demonstrate that observation noise does not impact the robot’s performance using our algorithm. This experiment draws the observation from the $3\times 3$ region around the true position with a truncated Gaussian distribution. The value shown in the figures is at the center of the $3\times 3$ region. The results indicate that, despite increasing observation noise, the success rates remain relatively consistent across different levels. A similar trend is observed for the number of localizations, suggesting that our algorithm maintains robust performance in the presence of observation noise.
Figure 11: (a) Number of Localizations vs. Different Threshold Levels. (b) Success Rate vs. Different Threshold Levels
@@ V-F Effect of Risk Constraint
Figures 11(a) and 11(b) depict the number of localizations and success rates under varying levels of the risk threshold. The results demonstrate that both metrics are influenced by changes in the risk threshold. As the threshold increases, the robot adopts more conservative movement strategies to minimize risk, resulting in a higher frequency of localization and fewer failures. Conversely, with a lower risk threshold, the robot engages in more aggressive actions, leading to increased uncertainty and a higher likelihood of failure. Consequently, the success rate diminishes as the risk constraint becomes more lenient.
VI Conclusion
We developed a novel active localization approach termed RiskRL, which combines a chance-constrained planner with a particle filter (PF). Our approach aims to minimize localization actions while not exceeding failure probabilities. The chance-constrained planner determines when the robot moves or localizes and was implemented using constrained reinforcement learning (RL). The PF maintains the robot’s belief by processing noisy 2D pose observations from the environment and motion commands. We also use the PF’s belief to compute the observation for our RL planner. The current approach succeeds our prior work [2], which employed an algorithm that has a slower inference time and requires well-defined transition and observation noise models, making it unusable in real-time, real-world scenarios. Our results revealed three key findings. First, RiskRL consistently outperformed the baseline by at least 13%, while maintaining a similar number of localizations. This demonstrates the robot’s ability to optimize the timing of localization actions, achieving higher rewards. Second, the robot dynamically adjusts the frequency of localizations, showcasing its adaptability to varying scenarios and environmental conditions. Finally, through bi-criteria optimization, RiskRL effectively controls risk levels while maximizing performance, ensuring the robot operates safely. Notably, RiskRL demonstrates zero-shot transfer capabilities, handling new environments without retraining, underscoring its potential for real-world deployment.
We believe that our models hold the potential for further optimization to achieve improved results. One promising direction is the implementation of Evolving Rewards [28] to identify a more effective set of reward values. Additionally, we aim to extend the current formulation to encompass both continuous state and action spaces, with the expectation that this will facilitate deployment in real-world scenarios.

### v2

@@ V-C Comparing with Baseline
We compared all algorithms in terms of the number of localize actions and success rates (that is, reaching the goal). The SP, TP, and CC-POMCP results represent the average performance of each policy type. In our experiments, we set $c=0.4$ for general training, and the policy successfully achieves this threshold. Even though in the train, RiskRL had similar success rates to (M2xL) and (M3xL) while still having a relatively high number of localizations, our experiments show that our RiskRL planner generalized very well to the harder chesapeake and st_thomas environments (Figure 7) as it was able to achieve a success rate nearly double of all of the other planners while keeping the number of localizations only marginally higher than M3xL. In the chesapeake environment, RiskRL has a 57% success rate while the next highest planner has a 31% success rate while averaging 19.7 localizations. In the st_thomas environment, RiskRL has a 61% success rate while the next highest planner has a 29% success rate while averaging 22.4 localizations.
It is interesting to note that even though BaseRL averaged very few localizations in the chesapeake and st_thomas environments (4.8 and 2.6, respectively), it was able to achieve a success rate on par with M2xL.
Compared to CC-POMCP (Figure 6), BaseRL performs similarly with a slightly higher success rate and number of localizations in the train environment. However, in terms of inference time, CC-POMCP takes around 20 seconds to choose one action, while BaseRL and RiskRL only take a maximum of 20 milliseconds.
Figure 6: The training graphs for train, showing the number of localizations (top) and success rates (bottom) and $c$ set to 0.4.
Figure 7: This figure depicts the number of localizations (top) and success rates (bottom) in each environment.
@@ V-D Analysis of RL Policies
Figure 8: This figure shows the probability of localizing based on the agent’s estimated location. The yellow block denotes the start position, and the green block denotes the goal position. Note for the chesapeake and st_thomas environments, we only show a smaller segment of the environment. We ran each algorithm on each environment with a predefined path 250 times.
Figure 8 shows where each RL planner tends to localize based on the belief’s mean. In general, BaseRL localizes without considering the success constraint, leading the planner to localize more near obstacles or around turns where the percentage of particles collided is high. In contrast, RiskRL distributes localizations more uniformly along the path with higher concentrations near obstacles, leading to a more robust policy with a higher success rate.
@@ V-E Effect of Transition and Observation Noise
Figure 9: This figure shows the localization counts (top) and success rates (bottom) at different transition noises in train.
Varying Transition Noise. Figures 9 illustrate the number of localizations and success rates for RiskRL as we vary the transition noise and use no observation noise. Here, the allowable probability of failure was set to $c=0.4$. Unsurprisingly, the localization count increases as the transition noise increases. More crucially, we observe that RiskRL respects the constraint even at the highest transition noise, reaffirming the correctness of the algorithm.
Figure 10: This figure shows the localization counts (top) and success rates (bottom) for different observation noises, where each percentage is the probability of observing the true position.
Varying Observation Noise. We now study how the observation noise affects the performance of RiskRL with the default transition noise. Figures 10 demonstrate that observation noise does not impact the agent’s performance. The results indicate that, despite increasing observation noise, the success rates remain relatively consistent across different levels. A similar trend is observed for the number of localizations, suggesting that our algorithm maintains robust performance in the presence of observation noise.
Figure 11: This figure shows the localization counts (top) and success rates (bottom) for varying risk thresholds $c$.
@@ V-F Effect of Risk Constraint
Figures 11 demonstrate that the risk threshold $c$ influences the localization counts and success rates. As $c$ increases, the agent becomes more risk-averse, resulting in a higher frequency of localization and fewer failures. Conversely, with a lower risk threshold, the agent takes more risks, leading to increased uncertainty and a higher likelihood of failure. Consequently, the success rate diminishes as the risk constraint becomes more lenient.
VI Discussion
The proposed RiskRL framework balances localization frequency with risk constraints, demonstrating robust performance across different environments. While our approach generalizes well to unseen environments, its adaptability to significantly different domains (e.g., highly dynamic or adversarial settings) warrants further investigation. Additionally, although RiskRL reduces inference time compared to CC-POMCP, training remains computationally expensive, requiring extensive interaction with the environment.
VII Conclusion
We developed a novel active localization approach termed RiskRL, which combines a high-level, chance-constrained planner with a particle filter (PF). Our approach aims to minimize localization actions while not exceeding failure probabilities. The chance-constrained planner determines when the robot moves or localizes and was implemented using constrained reinforcement learning (RL). The PF maintains the robot’s belief, processing noisy motion commands and 2D pose observations from the environment. We also use the PF’s belief to compute a 2D observation vector for the RL planner. The current approach succeeds our prior work [2], which employed an algorithm that has a slower inference time and requires well-defined transition and observation noise models, making it unusable in real-time, real-world scenarios. Our results revealed three key findings. First, RiskRL was able to achieve at least a 26% higher success rate compared to the baselines when deployed in unseen test environments (chesapeake and st_thomas). This demonstrates the robot’s ability to optimize when to localize, achieving higher rewards. Second, the robot dynamically adjusts its localization frequency, showcasing that it can adapt to different scenarios and environmental conditions. Finally, through bi-criteria optimization, RiskRL effectively controls risk levels while maximizing performance, ensuring the robot operates safely. Notably, RiskRL demonstrates zero-shot transfer capabilities, handling new environments without retraining, underscoring its potential for real-world deployment.
Our future plans include further optimization to improve results (for example, by implementing Evolving Rewards [28]), exploring continuous state and action spaces, and evaluating in more realistic environments.